OpenStudy (anonymous):

An arithmetic sequence is shown below. You are not given many terms of this sequence, but you do have enough to fill in the missing number. What number goes in the blank? 12, _, 54, ... Someone show me how to slove this please . Gonna be asked many more and don't know how to work it out .

OpenStudy (anonymous):

the missing number is 30, the difference is 21 12+21=33 30+21=54

OpenStudy (anonymous):

but there is a certain equation to solve problems like this .

OpenStudy (noseboy908):

You're correct. It's t(n)=t1+d(n-1), where t1 is the first term, d is how much is added to a term to get the next one, and n is the number of the term. In this case, twelve is the first term, and 54 is the 3rd term. Thus, you can fill in t1 and n, as those are given. You can also fill in the left side, because you're given that t(n) there, and you know that that represents the t value for the highest number given. Make sense?

OpenStudy (anonymous):

Uhhh... nopee lol

OpenStudy (noseboy908):

Ok, let me put it this way

OpenStudy (anonymous):

21 :DDD

OpenStudy (noseboy908):

Correct! Therefore you get the "d", which tells you how much to add to a number to get the next one.

OpenStudy (anonymous):

Thank you ! :)

OpenStudy (noseboy908):

Welcome! This is exactly why I don't understand why people would just give answers. In the end it doesn't help in the long term.

OpenStudy (anonymous):

Exactly . I like to know what I'm doing so if a similar question pops up I don't ask again just do it myself with no problem .